OpenStudy (anonymous):

Which of the following represents ^5sqrt x^4

OpenStudy (anonymous):

rational exponents?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

\[\huge \sqrt[n]{x^m}=x^{\frac{m}{n}}\]

OpenStudy (anonymous):

I think it is x^4/5

OpenStudy (anonymous):

the numerator is the power, the denominator is the root and vice versa

OpenStudy (anonymous):

you are right

OpenStudy (anonymous):

Ok thank